# Branch actions and the structure lattice

### v1

@@ Abstract.
J. S. Wilson proved in 1971 an isomorphism between the structural lattice associated to a group belonging to his second class of groups with every proper quotient finite and the Boolean algebra of clopen subsets of Cantor’s ternary set. In this paper we generalize this isomorphism to the class of branch groups. Moreover, we show that for every faithful branch action of a group $G$ on a spherically homogeneous rooted tree $T$ there is a canonical $G$-equivariant isomorphism between the Boolean algebra associated with the structure lattice of $G$ and the Boolean algebra of clopen subsets of the boundary of $T$.
@@
The first author is supported by the Spanish Government, grant PID2020-117281GB-I00, partly with FEDER funds. The first author also acknowledges support from the Walter Gyllenberg Foundation from the Royal Physiographic Society of Lund.
@@
Let $G\leq\mathrm{Aut}~{}T$ be a just-finite branch group. Then its structure lattice $\mathcal{L}(G)$ is isomorphic to the lattice of clopen subsets of Cantor’s ternary set.
@@
Our goal is to give a complete description of the structure lattice of a branch group. Indeed, we obtain a $G$-equivariant and explicit canonical isomorphism of Boolean algebras characterising the structure lattice of a branch group $G$:
@@
We use the exponential notation for the right action of a group on a tree and on its boundary. We write $H\leq_{f}$ if $H$ is a finite-index subgroup of $G$.
@@ 2.1. Boolean algebras
$0\vee 0=0\wedge 0=0\wedge 1=1\wedge 0=0,\quad 1\vee 1=1\wedge 1=1\wedge 0=0\wedge 1=1$
@@
$a\wedge 0=0,\quad a\wedge 1=a,\quad a\vee 0=a\quad\text{and}\quad a\vee 1=a,$
@@
Note that by definition $\neg 0=1$ and $\neg 1=0$.
@@ 2.2. Homomorphisms of Boolean algebras
Definition 2.6 (Homomorphism of Boolean algebra).
Let $A$ and $B$ be two Boolean algebras. We say a map $f:A\to B$ is a homomorphism of Boolean algebras if it preserves meets, joins, and both the maximal and the minimal elements $0,1$ respectively. In other words
$f(a\wedge b)=f(a)\wedge f(b),\quad f(a\vee b)=f(a)\vee f(b),\quad f(0)=0\quad\text{ and }\quad f(1)=1$
@@
First note that if $f$ is a homomorphism of boolean algebras then the set $I:=\{a\in A\mid f(a)=0\}$ is an ideal of $A$ as it is the kernel of the homomorphism $f$. Furthermore it is maximal by Lemma 2.9 as for every $a\in A$ either $f(a)=0$ or $f(\neg a)=0$ but not both. Now let us assume that the set $I$ is a maximal ideal of $A$ and let us prove that $f$ is then a homomorphism. We only need to show that if $f(a)=1$ for $a\in A$, then $f(a\vee b)=1$ for any $b\in A$, as the other conditions follow directly from the axioms on the definition of an ideal and $I$ being maximal. Note that $\neg a\in I$ by Lemma 2.8, thus $\neg a\wedge\neg b$ is also in $I$. Hence
¬($\neg(\neg a\wedge\neg b)=a\vee b$
is not in $I$ by Lemma 2.8 and therefore $f(a\vee b)=1$ as wanted. ∎
@@ 2.3. Stone spaces
A Stone space is also widely known as a profinite space; see [12]. The following lemma tells us how the clopen subsets of a Stone space look like.
@@
Since $C$ is open it is an arbitrary union of open subsets in $\mathcal{U}$, i.e. $C=\bigcup_{i\in I}U_{i}$. Since $C$ is closed in $X$ and $X$ is a Stone space $C$ is compact, thus there exists a finite refinement $C=\bigcup_{i=1}^{n}U_{i}$. To obtain a disjoint union, we remove unnecessary terms from the finite decomposition $C=\bigcup_{i=1}^{n}U_{i}$ using the fact that $U_{i}\cap U_{j}$ is clopen and equal to either $U_{i},U_{j}$ or $\emptyset$. ∎
An important example of Stone spaces is the Stone space associated to a Boolean algebra. Given a Boolean algebra $A$ its Stone space is simply $\mathrm{Hom}(A)$, i.e. the space consisting of all Boolean algebra homomorphisms $A\to\mathbf{2}$, which is a Stone space for the topology of pointwise convergence of nets. This topology is no more than the Tychonoff’s topology in $\mathrm{Hom}(A)$ when seen as the cartesian product $\mathbf{2}^{A}$.
@@
Let $X$ be a topological space such that $\mathrm{Bool}(X)\neq\mathbf{2}$. Then for every homomorphism of Boolean algebras $f:\mathrm{Bool}(X)\to A$ there exists some proper clopen $C$ such that $f(C)=1$.
@@
Since $\mathrm{Bool}(X)\neq\mathbf{2}$ there must exist a proper clopen subset $\emptyset\neq U\subset X$. Thus, its complement $\neg U$ is also a proper clopen subset of $X$. Then we have
@@
The map $\widetilde{f}$ is bijective and continuous (with continuous inverse) by definition, thus $\widetilde{f}\in\mathrm{Homeo}~{}X$. The map $F^{\prime}:\mathrm{Homeo}~{}X\to\mathrm{Aut}(\mathrm{Bool}(X))$ given by $f\mapsto f^{\prime}$ is a group homomorphism and similarly the map $\widetilde{F}:\mathrm{Aut}(\mathrm{Bool}(X))\to\mathrm{Homeo}~{}X$ given by $f\mapsto\widetilde{f}$ is a group homomorphism. Finally $\widetilde{F}\circ F^{\prime}=F^{\prime}\circ\widetilde{F}=\mathrm{id}$ as
@@
for any $x\in X$ and
$(\widetilde{f})^{\prime}(C)=f(C)$
for any clopen subset $C\subseteq X$. Hence $\mathrm{Homeo}~{}X$ is isomorphic to $\mathrm{Aut}(\mathrm{Bool}(X))$. ∎
@@ 3.1. Spherically homogeneous rooted trees and their boundaries
A spherically homogeneous rooted tree $T$ is a rooted tree such that every vertex in $T$ at the same distance from the root has the same degree. For every $n\geq 1$, the vertices at distance $n$ from the root constitute the $n$th level of $T$, denoted $\mathcal{L}_{n}$.
The tree $T$ can be identified with the set of finite words $\prod_{n\geq 1}X_{n}$, where the cardinality of each set $X_{n}$ coincides with the degree of the vertices at level $n$ for $n\geq 1$. Note that under this identification the root is simply the empty word. Also this identification induces a graded lexicographical order in $T$, by fixing a lexicographical order in $X_{n}$ for every $n\geq 1$.
Two vertices $u,v\in T$ are said to be incomparable if the unique paths from $u$ and $v$ to the root of $T$ intersect only at the root.
@@
The boundary $\partial T$ is a Stone space with respect to the compact topology generated by the cone sets $\{C_{v}\}_{v\in T}$. Cone sets are clopen sets in this topology and they satisfy the assumptions in Lemma 2.11. Thus any clopen in $\partial T$ may be represented as a finite disjoint union of cone sets. For a vertex $v\in T$ and a clopen $C\subseteq\partial T$ we write $v\in C$ if $C_{v}\subseteq C$.
@@ 3.2. The Boolean algebra of clopen sets of $\partial T$
Let us restrict the homomorphism $f$ to the subgraph $\{C_{v}\}_{v\in T}\subset\mathrm{Bool}(\partial T)$, which is isomorphic as a graph to $T$. The set $T$ is countable and thus the subset $S:=\{C_{v}\mid f(C_{v})=1\text{ and }v\in T\}$ is also countable, so let us write $S=\{C_{i}\}_{i\geq 1}$. By Lemmata 2.11 and 2.12 there exists at least one $v\in T$ such that $f(C_{v})=1$. Therefore for any $w$ above $v$ we must have $f(C_{w})=1$, as otherwise we would get $f(C_{v})=f(C_{v})\wedge f(C_{w})=0$. Thus the set $S$ is infinite. Furthermore since $f$ is a homomorphism, for any $C,D\in S$ we have
@@
However, two cone sets are either disjoint or contained one in the other. Thus since $C\cap D=\emptyset$ would imply $f(C\wedge D)=0$, we must have either $C\subseteq D$ or $D\subseteq C$ for any pair $C,D\in S$. This implies that the countable set $S$ is completely linearly ordered. Therefore $\gamma:=\bigcap_{C\in S}C$ is a well-defined uniquely determined element in the boundary $\partial T$. Note that $C_{v}\in S$ if and only if $\gamma\in C_{v}$.
Let us conclude by showing $f=\varphi_{\gamma}$. Let $A\in\mathrm{Bool}(\partial T)$. If $A$ contains $\gamma$, then $A$ must contain a cone set $C\in S$ by Lemma 2.11 and $f(A)=f(A\vee C)=1$. On the other hand, if $A$ does not contain $\gamma$, then $\neg A$ must contain $\gamma$ and we get $f(A)=\neg f(\neg A)=\neg 1=0$. Thus $f=\varphi_{\gamma}$. ∎
@@ 4.1. Branch groups
Let $T$ be a spherically homogeneous rooted tree and $\mathrm{Aut}~{}T$ its group of graph automorphisms. Let us fix a subgroup $G\leq\mathrm{Aut}~{}T$. We say $G$ is level-transitive if its action on every level of $T$ is transitive. We define the support of $G$, denoted $\mathrm{Supp}(G)$, via
@@
For a vertex $v\in T$, we write $\mathrm{st}_{G}(v)$ for the stabilizer of $v$ in $G$. We further define the rigid vertex stabilizer $\mathrm{rist}_{G}(v)$ of $v$ as the subgroup of $G$ consisting of elements whose support is contained in the subtree rooted at $v$. We get from definition
@@ 4.2. The structure lattice
$[H]\vee[K]:=[H\cap K]\quad\text{and}\quad[H]\wedge[K]:=[\langle H,K\rangle],$
@@
    it admits the distinguihsed elements $0:=[\{1\}]$ and $1:=[G]$;
@@ 4.3. Proof of A
We conclude the section by proving A. We shall fix a branch action $\rho:G\to\mathrm{Aut}~{}T$ and write $G\leq\mathrm{Aut}~{}T$ for the remainder of the section. First we show that the support of a subgroup with finite index normalizer is clopen:
@@
Let $G\leq\mathrm{Aut}~{}T$ a branch group and let $H\leq G$ be a subgroup with finitely many conjugates in $G$. Then $\mathrm{Supp}(H)$ is clopen.
@@
For any vertex $v\in T$ either $v$ is fixed by $H$ or $C_{v}\subseteq\mathrm{Supp}(H)$. Thus since $T$ is countable we can decompose $\mathrm{Supp}(H)$ as the disjoint countable union
@@
for some subset $V\subset T$. In particular $\mathrm{Supp}(H)$ is always open in $\partial T$. It remains to prove that $\mathrm{Supp}(H)$ is closed. We claim that the subset $V\subseteq T$ in Equation 4.1 is finite. This implies that $\mathrm{Supp}(H)$ is a finite union of clopen sets and thus it is clopen itself. Then let us prove our claim.
Assume by contradiction that $V$ is infinite. We shall assume that $V$ is ordered by the graded lexicographical order. Furthermore we may assume that $V$ does not contain any subset $W\subseteq V$ such that
@@
for some $v\in T$, as one may replace $W$ with $v$. Note that for any $g\in G$ we have
(4.3) $\mathrm{Supp}(H^{g})=\bigsqcup_{v\in V}C_{v^{g}}.$
We shall construct infinitely many distinct conjugates of $H$ which yields a contradiction. For that, it is enough to find an infinite sequence of levels $\{\ell_{n}\}_{n\geq 1}$ and of elements $\{g_{n}\}_{n\geq 1}\subseteq G$ such that $H^{g_{n+1}}$ and $H^{g_{n}}$ move the same vertices at level $\ell_{n}$ but move different ones at level $\ell_{n+1}$. We fix $\ell_{1}=1$ and $g_{1}=1$. We argue by induction on $n\geq 1$. Let $N_{n}\geq 1$ be such that $\mathrm{Rist}_{G}(\ell_{n})$ acts level-transitively on the subtrees rooted at level $\ell_{n}+N_{n}$, which is well-defined by [11, Lemma 2.5]. Since $V$ is infinite, there is $w_{n}\in V$ at some level $\ell_{n+1}$ such that $\ell_{n+1}>\ell_{n}+N_{n}$. Then, let us consider the unique vertices $u_{n}$ and $v_{n}$ above $w_{n}$ at levels $\ell_{n}$ and $\ell_{n}+N_{n}$ respectively. Consider $h_{n}\in\mathrm{rist}_{G}(u_{n})$ such that $w_{n}^{h_{n}}\notin V$, which exists by level-transitivity of $\mathrm{rist}_{G}(u_{n})$ on the subtree rooted at $v_{n}$ and the assumption in Equation 4.2. Let $g_{n+1}:=g_{n}h_{n}$. Then $H^{g_{n}}$ and $H^{g_{n+1}}$ move the same vertices at level $\ell_{n}$ but different ones at level $\ell_{n+1}$ level as wanted. ∎
@@
We need the following lemma in [1], which is stated more generally for weakly branch groups, where a weakly branch group $G$ is a level-transitve subgroup of $\mathrm{Aut}~{}T$ such that $\mathrm{Rist}_{G}(n)\neq 1$ for every $n\geq 1$.
@@
Let $G\leq\mathrm{Aut}~{}T$ be a weakly branch group and $H$ a $k$-subnormal subgroup of $G$. Then $H\geq\mathrm{rist}_{G}(v)^{(k)}$ for any $v\in\mathrm{Supp}(H)$.
@@
Now we show that $\mathrm{rist}_{G}(v)^{(k)}$ is subnormal in $G$ and that it has the same number of distinct conjugates as $H$. First $\mathrm{rist}_{G}(v)$ is subnormal in $G$ and thus $\mathrm{rist}_{G}(v)^{(k)}$ is subnormal in $G$ too as it is normal in $\mathrm{rist}_{G}(v)$. We know that $N_{G}(\mathrm{rist}_{G}(v))=\mathrm{st}_{G}(v)$, which is of finite index in $G$. Now, since $\mathrm{rist}_{G}(v)^{(k)}$ is characteristic in $\mathrm{rist}_{G}(v)$, it is normal in $\mathrm{st}_{G}(v)$ and therefore

### v2

@@ Abstract.
J. S. Wilson proved in 1971 an isomorphism between the structural lattice associated to a group belonging to his second class of groups with every proper quotient finite and the Boolean algebra of clopen subsets of Cantor’s ternary set. In this paper we generalize this isomorphism to the class of branch groups. Moreover, we show that for every faithful branch action of a group $G$ on a spherically homogeneous rooted tree $T$ there is a canonical $G$-equivariant isomorphism between the Boolean algebra associated to the structure lattice of $G$ and the Boolean algebra of clopen subsets of the boundary of $T$.
@@
The first author is supported by the Spanish Government, grant PID2020-117281GB-I00, partly with FEDER funds. The first author also acknowledges support from the Walter Gyllenberg Foundation from the Royal Physiographic Society of Lund. The second author is supported by Travel Support for Mathematicians, grant MP-TSM-00002045 from Simons Foundation.
@@
Let $G\leq\mathrm{Aut}~{}T$ be a just-infinite branch group. Then its structure lattice $\mathcal{L}(G)$ is isomorphic to the lattice of clopen subsets of Cantor’s ternary set.
@@
Our goal is to give a complete description of the structure lattice of a branch group. Indeed, we obtain a $G$-equivariant and explicit canonical isomorphism of Boolean algebras characterising the structure lattice of a (non-necessarily just-infinite) branch group $G$:
@@
We use the exponential notation for the right action of a group on a tree and on its boundary. We write $H\leq_{f}G$ if $H$ is a finite-index subgroup of $G$.
@@ 2.1. Boolean algebras
$0\vee 0=0\wedge 0=0\wedge 1=1\wedge 0=0,\quad 1\vee 1=1\wedge 1=1\vee 0=0\vee 1=1$
@@
$a\wedge 0=0,\quad a\wedge 1=a,\quad a\vee 0=a,\quad\text{and}\quad a\vee 1=1$
@@
Note that by definition $\neg 0=1$ and $\neg 1=0$ and recall that for every $a,b\in A$ we have de Morgan’s laws: $\neg(a\vee b)=\neg a\wedge\neg b$ and $\neg(a\wedge b)=\neg a\vee\neg b$.
@@ 2.2. Homomorphisms of Boolean algebras
Definition 2.6 (Homomorphism of Boolean algebras).
Let $A$ and $B$ be two Boolean algebras. We say a map $f:A\to B$ is a homomorphism of Boolean algebras if it preserves meets, joins, and the distinguished elements $0,1$ respectively. In other words
$f(a\wedge b)=f(a)\wedge f(b),\quad f(a\vee b)=f(a)\vee f(b),\quad f(0)=0,\quad\text{ and }\quad f(1)=1$
@@
A direct consequence of the above definition is that a homomorphism $f:A\to B$ preserves complements, i.e. $f(\neg a)=\neg f(a)$ for every $a\in A$.
@@
First note that if $f$ is a homomorphism of boolean algebras then the set $I:=\{a\in A\mid f(a)=0\}$ is an ideal of $A$ as it is the kernel of the homomorphism $f$. Furthermore it is maximal by Lemma 2.8 as for every $a\in A$ either $f(a)=0$ or $f(\neg a)=\neg f(a)=\neg 1=0$ but not both $f(a)=f(\neg a)=0$, as otherwise $f(1)=f(a\vee\neg a)=f(a)\vee f(\neg a)=0$. Now let us assume that the set $I$ is a maximal ideal of $A$ and let us prove that $f$ is then a homomorphism. We only need to show that if $a,b\notin I$ then $a\wedge b\notin I$ and that for every $a\in A$ if $b\notin I$ then $a\vee b\notin I$, as the other conditions follow directly from the axioms on the definition of an ideal and $I$ being maximal. If $a,b\notin I$ then $\neg(a\wedge b)=\neg a\vee\neg b\in I$ by Lemma 2.8 and thus $a\wedge b\notin I$ again by Lemma 2.8. Lastly if $a\in A$ and $b\notin I$ then $\neg(a\vee b)=\neg a\wedge\neg b\in I$ by Lemma 2.8 and hence $a\vee b\notin I$ also by Lemma 2.8. ∎
@@ 2.3. Stone spaces
A Stone space is also widely known as a profinite space; see [12]. The following lemma tells us how the clopen subsets of a Stone space look like:
@@
Since $C$ is open it is an arbitrary union of open subsets in $\mathcal{U}$, i.e. $C=\bigcup_{i\in I}U_{i}$. Since $C$ is closed in $X$ and $X$ is a Stone space $C$ is compact, thus there exists a finite refinement $C=\bigcup_{i=1}^{n}U_{i}$. To obtain a disjoint union, we remove unnecessary terms from the finite decomposition $C=\bigcup_{i=1}^{n}U_{i}$ using the fact that for $1\leq j<k\leq n$, the intersection $U_{j}\cap U_{k}$ is clopen and equal to either $U_{j},U_{k}$ or $\emptyset$. ∎
An important example of a Stone space is the Stone space associated to a Boolean algebra. Given a Boolean algebra $A$ its Stone space is simply $\mathrm{Hom}(A)$, i.e. the space consisting of all Boolean algebra homomorphisms $A\to\mathbf{2}$, which is a Stone space for the topology of pointwise convergence of nets. This topology is no more than the Tychonoff topology in $\mathrm{Hom}(A)$ when seen as the cartesian product $\mathbf{2}^{A}$.
@@
Let $X$ be a topological space such that $\mathrm{Bool}(X)\neq\mathbf{2}$. Then for every homomorphism of Boolean algebras $f:\mathrm{Bool}(X)\to\mathbf{2}$ there exists some non-empty proper clopen $C$ such that $f(C)=1$.
@@
Since $\mathrm{Bool}(X)\neq\mathbf{2}$ there must exist a proper clopen subset $\emptyset\neq U\subset X$. Thus, its complement $\neg U$ is also a non-empty proper clopen subset of $X$. Then we have
@@
The map $\widetilde{f}$ is bijective and continuous (with continuous inverse) by definition, thus $\widetilde{f}\in\mathrm{Homeo}~{}X$. The map $F^{\prime}:\mathrm{Homeo}~{}X\to\mathrm{Aut}(\mathrm{Bool}(X))$ given by $f\mapsto f^{\prime}$ is a group homomorphism and similarly the map $\widetilde{F}:\mathrm{Aut}(\mathrm{Bool}(X))\to\mathrm{Homeo}~{}X$ given by $f\mapsto\widetilde{f}$ is a group homomorphism. Finally $\widetilde{F}\circ F^{\prime}=\mathrm{id}_{\mathrm{Homeo~{}}X}$ and $F^{\prime}\circ\widetilde{F}=\mathrm{id}_{\mathrm{Aut(Bool}(X))}$ as
@@
for any $x\in X$ and $f\in\mathrm{Homeo}~{}X$, and
$(\widetilde{g})^{\prime}(C)=g(C)$
for any clopen subset $C\subseteq X$ and $g\in\mathrm{Aut}(\mathrm{Bool}(X))$. Hence $\mathrm{Homeo}~{}X$ is isomorphic to $\mathrm{Aut}(\mathrm{Bool}(X))$. ∎
@@ 3.1. Spherically homogeneous rooted trees and their boundaries
A spherically homogeneous rooted tree $T$ is a rooted tree such that every vertex in $T$ at the same distance from the root has the same degree. For every $n\geq 1$, the vertices at distance $n$ from the root constitute the $n$th level of $T$, denoted $\mathcal{L}_{n}$. For a vertex $v\in T$, the subtree rooted at $v$, denoted $T_{v}$, consists of the vertices in $T$ below $v$.
The tree $T$ can be identified with the set of finite words $\prod_{n\geq 1}X_{n}$, where the cardinality of each set $X_{n}$ coincides with the degree of the vertices at level $n-1$ for $n\geq 1$. Note that under this identification the root is simply the empty word. Also this identification induces a graded lexicographical order in $T$, by fixing a lexicographical order in $X_{n}$ for every $n\geq 1$.
Two vertices $u,v\in T$ are said to be incomparable if none is a descendant of the other.
@@
The boundary $\partial T$ is a Stone space with respect to the compact topology generated by the cone sets $\{C_{v}\}_{v\in T}$. Cone sets are clopen sets in this topology and they satisfy all the assumptions in Lemma 2.11. Thus any clopen in $\partial T$ may be represented as a finite disjoint union of cone sets. For a vertex $v\in T$ and a clopen $C\subseteq\partial T$ we write $v\in C$ if $C_{v}\subseteq C$.
@@ 3.2. The Boolean algebra of clopen sets of $\partial T$
Let us restrict the homomorphism $f$ to the subgraph $\{C_{v}\}_{v\in T}\subset\mathrm{Bool}(\partial T)$, which is isomorphic as a graph to $T$. By Lemmata 2.11 and 2.12 there exists a vertex $v\in T$ distinct from the root such that $f(C_{v})=1$. Again by Lemmata 2.11 and 2.12 there exists a vertex $v_{2}\in T_{v_{1}}\setminus\{v_{1}\}$ such that $f(C_{v_{2}})=1$. Applying this argument inductively shows that the set $S:=\{C_{v}\mid f(C_{v})=1\}$ is infinite. Furthermore since $f$ is a homomorphism, for any $C,D\in S$ we have
@@
However, two cone sets are either disjoint or contained one in the other. Thus since $C\cap D=\emptyset$ would imply $f(C\cap D)=0$, we must have either $C\subseteq D$ or $D\subseteq C$ for any pair $C,D\in S$. This implies that the countable set $S$ is completely linearly ordered. Therefore $\gamma:=\bigcap_{C\in S}C$ is a well-defined uniquely determined element in the boundary $\partial T$. Note that $C_{v}\in S$ if and only if $\gamma\in C_{v}$.
Let us conclude by showing $f=\varphi_{\gamma}$. Let $A\in\mathrm{Bool}(\partial T)$. If $A$ contains $\gamma$, then $A$ must contain a cone set $C\in S$ by Lemma 2.11 and $f(A)=f(A\cup C)=f(A)\vee f(C)=1$. On the other hand, if $A$ does not contain $\gamma$, then $\neg A$ must contain $\gamma$ and we get $f(A)=\neg f(\neg A)=\neg 1=0$. Thus $f=\varphi_{\gamma}$. ∎
@@ 4.1. Branch groups
Let $T$ be a spherically homogeneous rooted tree and $\mathrm{Aut}~{}T$ its group of graph automorphisms. Let us fix a subgroup $G\leq\mathrm{Aut}~{}T$. We say $G$ is level-transitive if its action on every level of $T$ is transitive. The group $G$ has a natural induced action on $\partial T$. We define the support of $G$, denoted $\mathrm{Supp}(G)$, via
@@
For a vertex $v\in T$, we write $\mathrm{st}_{G}(v)$ for the stabilizer of $v$ in $G$. We further define the rigid vertex stabilizer $\mathrm{rist}_{G}(v)$ of $v$ as the subgroup of $\mathrm{st}_{G}(v)$ consisting of elements whose support is contained in $C_{v}$. We get from definition that
@@ 4.2. The structure lattice
$[H]\wedge[K]:=[H\cap K]\quad\text{and}\quad[H]\vee[K]:=[\langle H,K\rangle],$
@@
    it admits the distinguished elements $0:=[\{1\}]$ and $1:=[G]$;
@@ 4.3. Proof of A
We conclude the section by proving A. We shall fix a branch action $\rho:G\to\mathrm{Aut}~{}T$ and write $G\leq\mathrm{Aut}~{}T$ for the remainder of the section. First we show that the support of a subgroup with finite index normalizer is clopen. For that we shall need the following lemma in [11]:
Lemma 4.2 (see [11, Lemma 2.5]).
Let $G\leq\mathrm{Aut}~{}T$ be a branch group. Then for every $n\geq 1$, there exists $N_{n}\geq 1$ such that $\mathrm{Rist}_{G}(n)$ acts level-transitively on the subtrees rooted at level $n+N_{n}$.
@@
Let $G\leq\mathrm{Aut}~{}T$ be a branch group and let $H\leq G$ be a subgroup with finitely many conjugates in $G$. Then $\mathrm{Supp}(H)$ is clopen.
@@
For any vertex $v\in T$ either $v$ is fixed by $H$ or $C_{v}\subseteq\mathrm{Supp}(H)$. Thus since $G\leq\mathrm{Aut}~{}T$, we can check which vertices are moved and which ones are fixed by $H$ at each level of $T$ and hence decompose $\mathrm{Supp}(H)$ as the disjoint countable union
@@
for some subset $V\subset T$ of pairwise incomparable vertices all moved by $H$. Note that in particular Equation 4.1 shows that $\mathrm{Supp}(H)$ is always open in $\partial T$. It remains to prove that $\mathrm{Supp}(H)$ is closed. We claim that the subset $V\subseteq T$ in Equation 4.1 is finite. This implies that $\mathrm{Supp}(H)$ is a finite union of clopen sets and thus it is clopen itself. Then let us prove our claim.
Assume by contradiction that $V$ is infinite. We shall assume that $V$ is ordered by graded lexicographical order. Furthermore we may assume that $V$ does not become finite if for any subset $W\subseteq V$ such that
@@
for some $v\in T$, one replaces $W$ with $v$ in $V$. Note that for any $g\in G$ we have
(4.3) $\mathrm{Supp}(H^{g})=\bigsqcup_{v\in V}C_{v^{g}}=\bigsqcup_{w\in V^{g}}C_{w},$
where $V^{g}$ consists of infinitely many pairwise incomparable vertices all moved by $H^{g}$.
We shall construct infinitely many distinct conjugates of $H$ which yields a contradiction. For that, it is enough to find an infinite sequence of levels $\{\ell_{n}\}_{n\geq 1}$ and of elements $\{g_{n}\}_{n\geq 1}\subseteq G$ such that for every $n\geq 1$ the subgroups $H^{g_{n+1}}$ and $H^{g_{n}}$ have the same action on level $\ell_{n}$ but they move different vertices at level $\ell_{n+1}$. We fix $\ell_{1}=1$ and $g_{1}=1$. We construct these two sequences by induction on $n\geq 1$. Let $N_{n}\geq 1$ be such that $\mathrm{Rist}_{G}(\ell_{n})$ acts level-transitively on the subtrees rooted at level $\ell_{n}+N_{n}$, which is well-defined by Lemma 4.2. Now, there is $w_{n}\in V^{g_{n}}$, i.e. $w_{n}$ is moved by $H^{g_{n}}$, at some level $\ell_{n+1}$ such that:
  1. (i)
    $\ell_{n+1}>\ell_{n}+N_{n};$
  2. (ii)
    if $v_{n}$ denotes the unique vertex at level $\ell_{n}+N_{n}$ above $w_{n}$ then there is a descendant $\widetilde{w}_{n}$ of $v_{n}$ at level $l_{n+1}$ fixed by $H^{g_{n}}$.
Note that condition (i) is guaranteed by $V^{g_{n}}$ being infinite while condition (ii) is guaranteed by Equation 4.3 and the assumption in Equation 4.2. Let $u_{n}$ be the unique vertex at level $\ell_{n}$ above both $w_{n}$ and $v_{n}$. Consider $h_{n}\in\mathrm{rist}_{G}(u_{n})$ such that $w_{n}^{h_{n}}=\widetilde{w}_{n}$, which exists by the level-transitivity of $\mathrm{rist}_{G}(u_{n})$ on the subtree rooted at $v_{n}$. Let $g_{n+1}:=g_{n}h_{n}$. Then $\widetilde{w}_{n}$ is moved by $H^{g_{n+1}}$ but it is fixed by $H^{g_{n}}$. Indeed, there exists $h\in H$ such that $w_{n}^{h^{g_{n}}}\neq w_{n}$, and thus
$\widetilde{w}_{n}^{h^{g_{n+1}}}=w_{n}^{h_{n}h_{n}^{-1}g_{n}^{-1}hg_{n}h_{n}}=w_{n}^{g_{n}^{-1}hg_{n}h_{n}}\neq w_{n}^{h_{n}}=\widetilde{w}_{n}.$
However, both subgroups $H^{g_{n+1}}$ and $H^{g_{n}}$ have the same action on level $\ell_{n}$, concluding the proof. ∎
@@
We need the following lemma in [1], which is stated more generally for weakly branch groups, where a weakly branch group $G$ is a level-transitive subgroup of $\mathrm{Aut}~{}T$ such that $\mathrm{Rist}_{G}(n)\neq 1$ for every $n\geq 1$.
@@
Let $G\leq\mathrm{Aut}~{}T$ be a weakly branch group and $H$ a $k$-subnormal subgroup of $G$. Then $H\geq\mathrm{rist}_{G}(v)^{(k)}$ for any $v\in T$ moved by $H$.
@@
Now we show that $\mathrm{rist}_{G}(v)^{(k)}$ is subnormal in $G$ and that it has the same number of distinct conjugates as $H$. First $\mathrm{rist}_{G}(v)$ is subnormal in $G$ as it is normal in the corresponding rigid level stabilizer, which is itself normal in $G$. Thus $\mathrm{rist}_{G}(v)^{(k)}$ is subnormal in $G$ too as it is itself normal in $\mathrm{rist}_{G}(v)$. We know that $N_{G}(\mathrm{rist}_{G}(v))=\mathrm{st}_{G}(v)$, which is of finite index in $G$. Now, since $\mathrm{rist}_{G}(v)^{(k)}$ is characteristic in $\mathrm{rist}_{G}(v)$, it is normal in $\mathrm{st}_{G}(v)$ and therefore